OpenStudy (vera_ewing):

Using the properties of exponents and radicals, design at least two different equivalent forms of x². You must show how each one can be simplified back to x² in two or more steps. Stretch your mind and get creative! Keep in mind that something too simple, like x • x would not be acceptable since it takes only one step to convert it to x².

OpenStudy (vera_ewing):

@satellite73

OpenStudy (anonymous):

how about \[\frac{x^{100}}{x^{98}}\]

OpenStudy (vera_ewing):

Does that have more than two steps?

OpenStudy (anonymous):

or \[x^7\times x^5\times x^{-10}\]

OpenStudy (vera_ewing):

What are the steps to simplify these?

OpenStudy (anonymous):

no that only has one the second one has more than one step add up the exponents